OpenStudy (mendicant_bias):

(PDE)(Sturm-Liouville) I'm solving an SLDE, and I'm not sure if I'm doing part of this right; I think I'm getting a trivial solution where I shouldn't be. Posted below in a second.

OpenStudy (mendicant_bias):

\[y''-\lambda y=0, \ \ \ y(0)=y(1)=0\]

OpenStudy (mendicant_bias):

What I've done: \[\text{Let $\lambda=-\alpha^2$, where $\alpha$ is a real constant. }\]\[y_h=c_1e^x+c_2e^-x, \ \ \ y'_h=c_1e^x-c_2e^{-x}\]\[y(0)=c_1+c_2=0.\]\[y(1)=c_1e+\frac{c^2}{e}=0.\]

OpenStudy (mendicant_bias):

\[(1) \ \ \ c_1+c_2=0\]\[(2) \ \ \ c_1-\frac{c^2}{e}=0\] (Note, the y(1) in the second post should be the same as the one in this one)

OpenStudy (mendicant_bias):

Substituting (1) into (2) in the form of c_1=-c_2 or vice versa, I just end up with \[c_1\bigg(e+\frac{1}{e}\bigg)=0\]

OpenStudy (mendicant_bias):

Is this wrong, or do I need to look for other values of lambda? (e.g. zero, +alpha^2) @Kainui

OpenStudy (kainui):

Are you sure you've written the problem correctly? I thought for a Sturm Liouville equation you have this form \[\Large \frac{d}{dx} \left( p(x) \frac{dy}{dx} \right) + q(x)y+ \lambda r(x) y =0\] with p(x), q(x), and r(x) all greater than 0. It appears as if your q(x)=0 and r(x)<0 but p(x)=1 is completely fine. Another thing, you have not really used the first derivative in your boundary conditions, which is why I thought it was odd you calculated it right before using your boundary conditions, but didn't plug it into anything, so I'm not really sure what's going on there.

OpenStudy (kainui):

Maybe somehow that negative sign will get absorbed into the eigenvalue based on how you've done it, but still something seems odd since having your coefficients both come to 0 seems to be either wrong or very simple solution. =P

OpenStudy (mendicant_bias):

That does have the general form of an SLDE if q(x)=0, p(x)=1, and r(x)=1. It's silly, but yeah, it does it the general form. Hmm.

OpenStudy (kainui):

Sure, but I'm saying q(x) needs to be larger than 0, I guess I need to review this method.

OpenStudy (mendicant_bias):

Sure thing, I'm reading it directly from the book and what you've said makes sense to me, too, but when I've brought up this exact same SLDE in a different context on OS and asked why this fit the general form, someone explained it to me as fitting the form as such. But yeah, I don't know what I'm doing, lol.

OpenStudy (mendicant_bias):

Brb, riding bike to different area, ~15 mins.

OpenStudy (dan815):

ya i do remember having to set some part of it to 0 so that it looks nice

OpenStudy (dan815):

i dontt thnk its q(x) though

OpenStudy (dan815):

in your case it is since that part doesnt exist